# Paying and Persuading

### v4

@@ 1 Introduction
Consider the problem faced by a principal who wishes to induce an agent to take an ex-ante undesirable action. Standard economic theory suggests two ways to motivate such an agent: directly compensate them for taking such an action with a monetary transfer (paying), or manipulate their information environment so that the action becomes ex-ante desirable (persuading). In settings where only one such tool is available to the principal, [^1] In moral hazard or incomplete contracting for the former, or information design for the latter. much is known about the optimal way to pay or persuade, and economists have derived various solutions in increasingly sophisticated models of both, separately.
@@
Despite this prima facie dichotomy between paying and persuading, there are new economic settings where a principal jointly controls information and monetary transfers. For example, in addition to varying payment, Uber strategically varies the amount of information drivers can see about recommended trips in order to motivate their drivers to accept rides that they may find ex-post undesirable. [^2] Uber writes “The type of approximate pickup and dropoff location shown to drivers varies based on address structures in your market, local regulations, and driver loyalty programs.” (Uber (2025)). Advertisers both run advertisements and email private offers that discount the good in order to incentivize purchases. Managers who privately observe the value or difficulty of a project may both strategically provide feedback to the worker and directly vary their compensation scheme in order to convince them to work when they otherwise might prefer to shirk.
@@
This paper studies dynamic contracting where Sender can both flexibly manipulate information and offer payments to Receiver. Formally, I consider a model where a principal (Sender) privately observes a Markovian state and then provides information and contingent transfers to an agent (Receiver) who chooses a perfectly monitored action in each period. Sender can commit to the entire path of transfers and information, conditioning payments and persuasion today based on anything in the past history. [^3] For example, Uber writes the algorithm for, and publicizes, their dynamic reward structure simultaneously, committing to rewarding drivers in the future for past performance. Information and transfers thus have an intertemporal component: they can be used immediately to change behavior today, but can also be used in the future as a reward for changes in behavior today as a form of dynamic compensation. Thus, when seeking to motivate Receiver, Sender must balance the use of information and transfers both within each period and across time.
@@
My main result is a general backloading theorem that sharply characterizes the tradeoff between optimal information and transfers in Sender’s dynamic contracting problem. Formally, Theorem 1 constructs an optimal contract where, if transfers are ever used, Sender has already committed to fully reveal the state at all continuation histories, i.e. committed to relinquishing their informational advantage at every point in the future. This characterization applies for any Sender or Receiver stage-game payoffs, discount rate, and law governing the Markov chain, and thus speaks generally to the way information and transfers trade off against one another in dynamic contracts. Theorem 1 can be interpreted as a sequencing argument: giving Receiver information tomorrow must always be weakly as efficient as simply paying Receiver today, and hence there is an optimal strategy where payments are made on the path of play only once information is exhausted as a dynamic incentive instrument.
The key step of the proof is a pullback lemma (Lemma 4) which establishes that the full-information experiment lies on the Pareto frontier (i.e. is an efficient way to provide some level of incentives). Providing more information has two opposing effects: it increases the total efficiency of the decision, increasing total surplus; however, it also increases Receiver’s relative share of the total surplus. Because Receiver ultimately takes the action, the second force can sometimes dominate the first, rendering (from Sender’s perspective) information a suboptimal way to provide incentives. Thus, in the absence of transfers, information provision can be a particularly expensive way to provide incentives. However, Lemma 4 shows that, so long as Sender has access to transfers, there is a way to counterbalance the second force so that the cost of providing more information in conjunction with transfers is always more efficient than simply paying Receiver. Thus, optimally paying and persuading tomorrow must be a weakly more efficient method of providing incentives than only paying today, implying payments can always be backloaded. Since Lemma 4 requires Sender to be able to jointly pay and persuade in the future, it highlights how transfers augment the cost of information revelation in dynamic persuasion.
@@
Following Theorem 1, I characterize several properties of actions and beliefs that hold uniformly under all optimal contracts for i.i.d. states. First, Propositions 3 and 4 identify testable conditions on the sets of actions and beliefs consistent with any optimal contract, regardless of whether transfers are fully backloaded. [^4] Recall Theorem 1 establishes one such contract where transfers are backloaded: Propositions 3 and 4 identify properties that hold uniformly across all optimal contracts. These restrictions are based on the ratio of Sender to Receiver payoffs when switching an action from $a$ to $a^{\prime}$ at a belief $\mu$, an object I call the effectiveness ratio. Second, Proposition 6 identifies conditions where all optimal contracts backload transfers, strengthening the conclusion of Theorem 1 under additional regularity conditions. These conditions ensure information and transfers are not always perfect substitutes, which is sufficient to guarantee that informational incentives always take strict precedence over transfers as a motivating tool.
@@
Next, I compute the value of paying and persuading in several specialized contexts. First, I consider the static joint design problem. Here, I show the value of joint information and transfers takes the form of an augmented concavification—which I call the $\mathcal{K}$-cavification. The $\mathcal{K}$-cavification requires Sender to compute their optimal transfer only at each of finitely many extremal beliefs—those where Receiver is indifferent between maximally many actions—and then linearly interpolates between the value at each of these beliefs to obtain the value of joint payments and persuasion. This iterative procedure dramatically simplifies the complexity of finding the value of jointly paying and persuading. I then show how this procedure gives simple conditions under which dynamic incentives are strictly optimal.
@@
Second, I bound the value of dynamically paying and persuading as discounting vanishes when the state transition process is ergodic. Sender’s value from the optimal dynamic contract must be bounded from below by their best payoff from all static couplings over states and actions that satisfy a Receiver incentive compatibility constraint evaluated at the ergodic distribution of the transition process. Such a characterization implies simple conditions under which Sender’s dynamic value strictly outperforms their one-shot persuasion value in the static game, and thus implies conditions where dynamic incentivization benefits Sender. It also allows me to comment on the literature on Markovian persuasion, and show that when Receiver is not myopic, the canonical upper bound established in the literature instead becomes a lower bound (Proposition 9).
@@
Finally, I explicitly characterize the optimal contract in an example inspired by rideshare design. Uber (Sender) must repeatedly incentivize a driver (Receiver) to accept multiple rides in each period. [^5] My stage game nests as a special case the leading binary-action binary-state example in the persuasion literature. I show that when the value of rides in each period is independently drawn and additively separable for both players, the optimal contract is a tiered loyalty program: Sender chooses the static optimal information structure until a random promotion time, after which Sender reveals the state and simply pays Receiver. This optimal contract features two stages—a first stage in which Uber counts the number of times drivers take rides they would prefer not to take and manipulates their informational environment (i.e. “Uber Blue”), and a second tier where Uber reveals the state and allows the driver to reject bad rides without penalty (i.e. “Uber Gold). My simple example thus fits with the publically available stylized description of Uber’s driver loyalty program.
@@ 1.1 Related Literature
Methodologically, I pull from the literature that characterizes the optimality of dynamic contracts by tracking the geometry of the Pareto frontier for incentives, an approach initiated by Spear and Srivastava (1987). Subsequently, it has been used in dynamic moral hazard problems by Lipnowski and Ramos (2020), and Guo and Hörner (2020), relational contracting by Li and Matouschek (2013), political economy by Yared (2010), Padró i Miquel and Yared (2012), and Acharya et al. (2024), and dynamic verification by Li and Libgober (2024). I contribute to this literature by using the properties of the Pareto frontier to characterize the tradeoff between two distinct tools Sender can access, and apply these methods to dynamic information design.
@@
Importantly, (and relative to the above literature), I suppose the principal can flexibly design the information structure, in the spirit of the Bayesian persuasion literature started by Rayo and Segal (2010) and Kamenica and Gentzkow (2011). A rich literature has since pushed the limits of static persuasion; see Bergemann and Morris (2019) for a survey. Yet work on dynamic persuasion is much sparser. Ely (2017), Renault et al. (2017) and Lehrer and Shaiderman (2025) assume a myopic Receiver and characterize conditions where greedy policies are optimal, while Ball (2023) solves the full dynamic persuasion problem but only for a linear-quadratic game with a Brownian state. In the absence of commitment, Golosov et al. (2014) and Renault et al. (2013) study repeated cheap talk games and characterize when full revelation is possible and the payoff set of the patient game, respectively.
@@
The literature on persuasion with transfers is similarly sparse. Li (2017) solves an example where Sender can pay and persuade in a static setting with binary actions, binary states, and state-independent payoffs. Further afield, Ravid et al. (2022), Terstiege and Wasser (2022), and Yamashita and Zhu (2023) study the joint design of information and transfers in auctions and screening, assuming the individual who persuades (Seller) is distinct from the one who pays (buyer). To the best of my knowledge, this paper is the first to consider a general model (static or dynamic) where the principal controls information and transfers, and the first to establish qualitative properties of the optimal mechanism. I also provide a new characterization of the value of dynamic persuasion (with payments), which nest the value of dynamic persuasion (by taking the cost of transfers to $\infty$).
@@
Also related is the literature on information design in optimal stopping problems (Ely and Szydlowski (2020), Orlov et al. (2020), Koh and Sanguanmoo (2024) and Koh et al. (2024)), which characterizes the optimal mechanism when Receiver takes an action only once and the state is perfectly persistent. Importantly, transfers are absent in this literature. Consequently the economic forces driving our models are quite different. Particularly related is Ely and Szydlowski (2020), who study optimal dynamic information design in a continuous time, work-or-shirk model when the transfer scheme is fixed and paid out only at the end; I show that when transfers can be endogenously combined with information in any flexible fashion, payments are paid out only after the initial (pathwise) period of pure information design.
@@
My geometric characterization of static persuasion with transfers (Proposition 8) relates to the literature on the geometry of persuasion. In particular, I build on work that characterizes the value of persuasion geometrically in Bayesian persuasion (Kamenica and Gentzkow (2011)), cheap talk with transparent motives (Lipnowski and Ravid (2020)), constrained persuasion (Doval and Skreta (2023)), and communication with flexible information acquisition (Barros (2025)); I identify the relevant object to be concavified in the presence of both transfers and persuasion.
Finally, I relate to a newer literature seeking to understand the impact of persuasion in repeated games, often with some sort of monetary transfer. Ortner et al. (2024) and Sugaya and Wolitzky (2025) study mediated games where a designer wishes to implement some optimal outcome among players who repeatedly play a pricing game (i.e. a first price auction or a Bertrand competition market game). Kolotilin and Li (2021) study a linear cheap-talk game and show Sender attains their best monotone contract when transfers are voluntary (i.e. relational). In contrast, I analyze a model where the person who persuades is also the one who pays, and give a backloading result when the principal can commit to the entire path of transfers.
@@ 2.1 The Stage Game
There are two players, Sender (S) and Receiver (R). Sender privately observes a state $\theta$ drawn from distribution $\mu_{0}\sim\Delta(\Theta)$; Receiver takes action $a\in A$. $A$ and $\Theta$ are both finite [^7] Except for the last step of the proof of Theorem 1 and Proposition 8, both $A$ and $\Theta$ need only be compact sets, modulo the appropriate measure-theoretic generalization. However, to invoke the strong topology (see Appendix B: The Strong Topology on Strategies ) and compute extremal beliefs, respectively, $A$ and $\Theta$ must be finite..
Sender can design a contract to motivate Receiver in one of two ways. First, they can transmit information, modeled as a choice of signal structure $\mathcal{S}:\Theta\to\Delta(S)$ from states into signals, where the signal space $S$ is a subset of some standard Borel space. Second, they can commit to a transfer rule, $t:\Theta\times S\times A\to\mathbb{R}_{+}$ which specifies a nonnegative [^8] In the absence of a nonnegativity restriction, i.e. a limited liability constraint, Sender can obtain first best by “informationally” selling the firm to the agent—offering full information for a payment equal to their value of information, then paying them to take the surplus maximizing action. payment for each pair of states, signals, and actions. Let $\mathcal{T}$ denote the set of all transfer rules. Having seen $(\mathcal{S},t)$ and realized signal $s\in S$, the agent forms beliefs and takes an action $a$ maximizing their expected transfer-augmented utility
@@
Let $v:A\times\Theta\to\mathbb{R}$ be Sender’s payoff. Following the approach of Kamenica and Gentzkow (2011) (modified for transfers), I represent a tuple $(S,t,\bar{a}(\cdot))$ where $\bar{a}(\cdot)$ best responds to $(\mathcal{S},t)$ as a tuple $(\rho,t,\bar{a}(\cdot))$ where $\rho\in\Delta_{\mu_{0}}(\Delta(\Theta))$ is a $\mu_{0}$-Bayes plausible distribution of posterior beliefs and
@@
An immediate simplification allows the transfer rule to be written only as a function of $(s,a)$, integrating out the state. This is because incentive compatibility of any recommended action depends only on the expected transfer given any signal, not the specific state-dependent ex-post payment.
@@
Fix any $(\mathcal{S},t,\bar{a})$. There is an outcome equivalent $(\mathcal{S}^{\prime},t,\bar{a})$ where $t(\cdot,s,a)$ is constant in $\theta$.
@@
Define $t(\theta,s,a)=\mathbb{E}_{\theta|s}[t(\theta,s,a)]$. Clearly, this induces the same expected transfers and joint distribution of beliefs and actions, maintaining incentive compatibility. ∎
A straightforward modification extends the lemma into the dynamic game described in the next subsection. I omit the dynamic generalization for expositional succinctness. In principle, a revelation principle also holds (see Kamenica and Gentzkow (2011)) where it is without loss of generality for $S=A$. However, it turns out that it will be useful in the dynamic case for the signal space to be $S=\Delta(\Theta)\times A$, i.e. for Sender to recommend a profile of actions at each belief.

### v5

@@ Abstract
Keywords: Information Design, Dynamic Mechanisms, Backloading.
JEL Codes: D82, D83, D86.
@@ 1 Introduction
Consider the problem faced by a principal who wishes to induce an agent to take an ex-ante undesirable action. Standard economic theory suggests two ways to motivate such an agent: directly compensate them for taking such an action with a monetary transfer (paying), or manipulate their information environment so that the action becomes interim desirable (persuading). In settings where only one such tool is available to the principal [^1] In moral hazard or incomplete contracting for the former, or information design for the latter., much is known about the optimal way to pay or persuade, and economists have derived various solutions in increasingly sophisticated models of both, separately.
@@
Despite this prima facie dichotomy between paying and persuading, there are new economic settings where a principal jointly controls information and monetary transfers. For example, in addition to varying payments, Uber strategically varies the amount of information drivers can see about recommended trips in order to motivate their drivers to accept rides that they may find ex-post undesirable. [^2] Uber writes “The type of approximate pickup and dropoff location shown to drivers varies based on address structures in your market, local regulations, and driver loyalty programs.” (Uber (2025)). Advertisers both run advertisements and email private offers that discount the good in order to incentivize purchases. Managers who privately observe the value or difficulty of a project may both strategically provide feedback to the worker and directly vary their compensation scheme in order to convince them to work when they otherwise might prefer to shirk.
@@
This paper studies dynamic contracting where Sender can both flexibly manipulate information and offer payments to Receiver. Formally, I consider a model where a principal (Sender) commits to (1) provide information about a Markovian state and (2) a contingent transfer rule prior to an agent (Receiver) choosing a perfectly monitored action in each period. Sender can commit to the entire path of transfers and information, conditioning payments and persuasion today based on anything in the past history. [^3] For example, Uber writes the algorithm for, and publicizes, their dynamic reward structure simultaneously, committing to rewarding drivers in the future for past performance. Information and transfers thus have an intertemporal component: while payments can occur immediately to change behavior today, they can also be leveraged in the future as a dynamic reward. Thus, when seeking to motivate Receiver, Sender must balance the use of information and transfers both within each period and across time.
@@
My main result is a striking backloading theorem that characterizes the tradeoff between optimal information and transfers in Sender’s dynamic contracting problem. Formally, Theorem 1 constructs an optimal contract where, if transfers are ever used, Sender has already committed to fully reveal the state at all continuation histories, i.e. committed to relinquishing their informational advantage at every point in the future. This characterization applies for any Sender or Receiver stage-game payoffs, discount rate, and law governing the Markov chain, and thus speaks generally to the way information and transfers trade off against one another in dynamic contracts. Theorem 1 can be interpreted as a sequencing argument: jointly paying and persuading Receiver tomorrow must always be weakly as efficient as only paying Receiver today, and hence there is an optimal contract where payments are made on the path of play only once information is exhausted as a dynamic incentive instrument.
The proof hinges on a pullback argument (Lemma 4), which establishes that the full-information experiment lies on the Pareto frontier. [^4] The Pareto frontier is the graph of Sender’s value function $V(u)$, i.e. their value from optimally providing utility $u$ to Receiver. This result depends on a careful interaction between information and transfers. In particular, it must show that transfers alone are not a cheaper way to provide incentives when compared to fully revealing the state. A-priori, one might worry that disclosure might be inefficient because Receiver would respond to new information in ways particularly unfavorable to Sender. Lemma 4 ameliorates this concern by showing that a joint payment-and-persuasion scheme must be at least as efficient as using payments alone. The key idea is simple: Sender can, after revealing the state, always pay Receiver to take any action that Sender wants at a constant efficiency cost. Thus additional information revelation cannot impose a higher utility cost than simply paying for the action Receiver would have taken in the absence of additional information. Consequently, combining future payments with persuasion is weakly more efficient than making an immediate payment, creating incentives to backload transfers. This illustrates how transfers fundamentally change the opportunity cost of information revelation in dynamic persuasion.
@@
Following Theorem 1, I characterize several properties of actions and beliefs that hold uniformly under all optimal contracts for i.i.d. states. First, Propositions 3 and 4 identify testable conditions on the sets of actions and beliefs consistent with any optimal contract, regardless of whether transfers are fully backloaded. [^5] Recall Theorem 1 establishes one such contract where transfers are backloaded: Propositions 3 and 4 identify properties that hold uniformly across all optimal contracts. These restrictions are based on the ratio of Sender to Receiver payoffs when switching an action from $a$ to $a^{\prime}$ at a belief $\mu$, an object I call the effectiveness ratio. Second, Propositions 6 and 7 identify conditions where all optimal contracts backload transfers, strengthening the conclusion of Theorem 1 under additional regularity conditions. These conditions ensure information and transfers are not always perfect substitutes, which is sufficient to guarantee that informational incentives always take strict precedence over transfers as a motivating tool. Finally, Proposition 8 shows eveyr optimal contract backloads transfers whenever transfers are sufficiently expensive.
@@
Next, I compute the value of paying and persuading in several specialized contexts. First, I consider the static joint design problem. Here, I show the value of joint information and transfers takes the form of an augmented concavification—which I call the $\mathcal{K}$-cavification (Proposition 9). The $\mathcal{K}$-cavification requires Sender to compute their optimal transfer only at each of finitely many extremal beliefs—those where Receiver is indifferent between maximally many actions—and then linearly interpolates between the value at each of these beliefs to obtain the value of joint payments and persuasion. This iterative procedure dramatically simplifies the complexity of finding the value of jointly paying and persuading and gives simple conditions under which dynamic incentives are strictly optimal in the general model.
@@
Second, I bound the value of dynamically paying and persuading as players become very patient when the state transition process is ergodic. Sender’s value from the optimal dynamic contract must be bounded from below by their best payoff from all static couplings over states and actions that satisfy a Receiver individual rationality constraint evaluated at the ergodic distribution of the transition process. Such a characterization implies simple conditions under which Sender’s dynamic value strictly outperforms their one-shot persuasion value in the static game, and thus implies conditions where dynamic incentivization benefits Sender. It also allows me to comment on the literature on Markovian persuasion, and show that when Receiver is not myopic, the canonical upper bound established in the literature instead becomes a lower bound (Proposition 10).
@@
Finally, I explicitly characterize the optimal contract in a rideshare-inspired example. Uber (Sender) repeatedly incentivizes a driver (Receiver) to accept multiple rides in each period. [^6] The stage game nests the canonical binary-state, binary-action example from the persuasion literature as a special case. When ride values are independently drawn across periods and utilities are additively separable, the optimal contract takes the form of a tiered loyalty program: In each tier, Uber chooses the static optimal information structure until a random promotion time, after which they reveal some states and simply pay the Receiver. The optimal contract therefore features distinct, ranked phases; in each phase, Uber tracks acceptance behavior for some number of periods and strategically manipulates the driver’s informational environment. As the relationship matures, it moves up in rank, and Uber begins to reveal more information about ride quality. Moreover, they allow the driver to reject undesirable rides without penalty, relying directly on payments for incentives. This characterization is reminiscent of stylized, publically available facts regarding Uber’s loyalty program for their drivers.
@@ 1.1 Related Literature
Methodologically, I pull from the literature that characterizes the optimality of dynamic contracts by tracking the geometry of the Pareto frontier for incentives, an approach initiated by Spear and Srivastava (1987). Subsequently, it has been used in dynamic moral hazard problems by Lipnowski and Ramos (2020) and Guo and Hörner (2020), relational contracting by Li and Matouschek (2013), political economy by Yared (2010), Padró i Miquel and Yared (2012), and Acharya et al. (2024), and dynamic verification by Li and Libgober (2024). I contribute to this literature by using the properties of the Pareto frontier to characterize the tradeoff between two distinct tools Sender can access, and apply these methods to dynamic information design.
@@
Importantly, I suppose the principal can flexibly design the information structure, in the spirit of the Bayesian persuasion literature started by Rayo and Segal (2010) and Kamenica and Gentzkow (2011). A rich literature has since pushed the limits of static persuasion; see Bergemann and Morris (2019) for a survey. Yet work on dynamic persuasion is sparser. Ely (2017), Renault et al. (2017) and Lehrer and Shaiderman (2025) assume a myopic Receiver and characterize when greedy policies are optimal, while Ball (2023) solves the full dynamic persuasion problem but only for a linear-quadratic game with a Brownian state. Further afield, Ely et al. (2015) study optimal information provision when Receiver has preferences for noninstrumental information. Doval and Ely (2020) and Makris and Renou (2023) characterize the set of all possible outcomes attainable in a dynamic information design problem, but leave open the question of optimal information design. In the absence of commitment, Golosov et al. (2014) and Renault et al. (2013) study repeated cheap talk games and characterize when full revelation is possible and the payoff set of the patient game, respectively.
@@
The literature on persuasion with transfers is similarly sparse. Li (2017) solves an example where Sender can pay and persuade in a static setting with binary actions, binary states, and state-independent payoffs. Further afield, Ravid et al. (2022), Terstiege and Wasser (2022), and Yamashita and Zhu (2023) study the joint design of information and transfers in auctions and screening, assuming the individual who persuades (Seller) is distinct from the one who pays (Buyer). In the one-shot context, Doval and Skreta (2022) fully characterize the optimal joint paying-and-persuading model in a mechanism design context by recasting limited commitment on the designer’s part as a Bayes plausibility constraint. Finally, in the dynamic context Hörner and Skrzypacz (2016) consider a binary-state game where the persuader is paid transfers for revealing information, and show that multi-stage communication is always beneficial to the high-type agent. Relatedly, Madsen et al. (2026) study a static, reduced-form model where a principal trades off between quasilinear transfers and convex amenity tools, interpreting the convex cost as continuation utility in a dynamic interaction reminiscent of my setting. To the best of my knowledge, this paper is the first to embed transfers into a dynamic model of persuasion, and the first to establish qualitative properties of the optimal mechanism. I also provide a new characterization of the value of dynamic persuasion with payments, which nest the value of dynamic persuasion alone (by taking the cost of transfers to be arbitrarily large).
@@
Also related is the literature on information design in optimal stopping problems (Ely and Szydlowski (2020), Orlov et al. (2020), Koh and Sanguanmoo (2024) and Koh et al. (2024)), which characterizes the optimal mechanism when Receiver takes an action only once and the state is perfectly persistent. Importantly, flexible endogenous transfers are absent in this literature. Consequently the economic forces driving our models are quite different. Particularly related is Ely and Szydlowski (2020), who study optimal dynamic information design in a continuous time, work-or-shirk model when the transfer scheme is fixed and paid out only at the end; I show that when transfers can be endogenously combined with information in any flexible fashion, payments are paid out only after the initial (pathwise) period of pure information design.
@@
My geometric characterization of static persuasion with transfers (Proposition 9) contributes to the literature on the geometry of communication. I build on work that geometrically characterizes the value of communication in persuasion (Kamenica and Gentzkow (2011)), cheap talk with transparent motives (Lipnowski and Ravid (2020)), constrained persuasion (Doval and Skreta (2023)), and flexible information acquisition (Barros (2025)).
Finally, I relate to a newer literature on persuasion in repeated games, often with monetary transfers. Ortner et al. (2024) and Sugaya and Wolitzky (2026) study mediated games where a designer wishes to implement some optimal outcome among players who repeatedly play a pricing game (i.e. a first price auction or a Bertrand competition market game). Kolotilin and Li (2021) study a linear cheap-talk game and show Sender attains their best monotone contract when transfers are voluntary (i.e. relational). In contrast, I analyze a model where the person who persuades is also the one who pays, and give a backloading result when the principal can commit to the entire path of transfers.
@@ 2.1 The Stage Game
There are two players, Sender (S) and Receiver (R). Sender privately observes a state $\theta$ drawn from distribution $\mu_{0}\sim\Delta(\Theta)$; Receiver takes action $a\in A$. $A$ and $\Theta$ are both finite. [^8] Except for the last step of the proof of Theorem 1 and Proposition 9, both $A$ and $\Theta$ need only be compact sets, modulo the appropriate measure-theoretic generalization. However, to invoke the strong topology (see Appendix B: The Strong Topology on Contracts ) and compute extremal beliefs, respectively, $A$ and $\Theta$ must be finite.
Sender can design a contract to motivate Receiver in one of two ways. First, they can transmit information, modeled as a choice of signal structure $\mathcal{S}:\Theta\to\Delta(S)$ from states into signals, where the signal space $S$ is a subset of some standard Borel space. Second, they can commit to a transfer rule, $t:\Theta\times S\times A\to[0,C]$ which specifies a nonnegative [^9] In the absence of a nonnegativity restriction, i.e. a limited liability constraint, Sender can obtain first best by “informationally” selling the firm to the agent—offering full information for a payment equal to their value of information, then paying them to take the surplus maximizing action. bounded [^10] I allow $C$ to be arbitrarily large, but finite for technical reasons; in the dynamic model, this ensures transfers satisfy a standard transversality condition that is crucial for compactness of the contract space. payment for each pair of states, signals, and actions. Let $\mathcal{T}$ denote the set of all transfer rules. Having seen $(\mathcal{S},t)$ and realized signal $s\in S$, the agent forms beliefs and takes an action $a$ maximizing their expected transfer-augmented utility
@@
Let $v:A\times\Theta\to\mathbb{R}$ be Sender’s bernoulli payoff. Following the approach of Kamenica and Gentzkow (2011) (modified for transfers), I represent a tuple $(S,t,\bar{a}(\cdot))$ where $\bar{a}(\cdot)$ best responds to $(\mathcal{S},t)$ as a tuple $(\rho,t,\bar{a}(\cdot))$ where $\rho\in\Delta_{\mu_{0}}(\Delta(\Theta))$ is a $\mu_{0}$-Bayes plausible distribution of posterior beliefs and
@@
    Nature draws $\theta\sim\mu_{0}$, and then signal $S$ is realized according to lottery $\mathcal{S}(\theta)$.
  4. (4)
@@
An immediate simplification allows the transfers to be written only as a function of $(s,a)$, integrating out the state. This is because incentive compatibility of any recommended action depends only on the expected transfer given any signal, not the specific state-dependent ex-post payment. Say tuples $(\mathcal{S},t,a)$ and $(\mathcal{S}^{\prime},t^{\prime},a^{\prime})$ are outcome equivalent if they induce the same expected on-path transfers and joint distribution over beliefs and actions.
@@
Fix $(\mathcal{S},t,\bar{a})$. There exists outcome equivalent $(\mathcal{S}^{\prime},t^{\prime},\bar{a})$ where $t^{\prime}$ is constant in $\theta$.
@@
Define $t^{\prime}(s,a)=\mathbb{E}_{\theta|s}[t(\theta,s,a)]$. Clearly, this induces the same expected transfers and joint distribution of beliefs and actions, maintaining incentive compatibility. ∎
A straightforward modification extends the lemma into the dynamic game described in the next subsection. I omit the dynamic generalization for expositional succinctness. In principle, a revelation principle also holds (see Kamenica and Gentzkow (2011)) where it is without loss of generality for $S=A$. However, it turns out that it will be useful in the dynamic case for the signal space to be $S=\Delta(\Theta)\times A$, i.e. for Sender to both specify a posterior belief (a la the belief-based approach), and a recommended action (a la the obedience approach).